# sangkuriang: A pseudo-spectral Python library for Korteweg-de Vries soliton simulation

### v2

@@ Abstract
Korteweg-de Vries (KdV) equation serves as a foundational model in nonlinear wave physics, describing the balance between dispersive spreading and nonlinear steepening that gives rise to solitons. This article introduces sangkuriang, an open-source Python library for solving this equation using Fourier pseudo-spectral spatial discretization coupled with adaptive high-order time integration. The implementation leverages just-in-time (JIT) compilation for computational efficiency while maintaining accessibility for instructional purposes. Validation encompasses progressively complex scenarios including isolated soliton propagation, symmetric two-wave configurations, overtaking collisions between waves of differing amplitudes, and three-body interactions. Conservation of the classical invariants is monitored throughout, with deviations remaining small across all test cases. Measured soliton velocities conform closely to theoretical predictions based on the amplitude-velocity relationship characteristic of integrable systems. Complementary diagnostics drawn from information theory and recurrence analysis confirm that computed solutions preserve the regular phase-space structure expected for completely integrable dynamics. The solver outputs data in standard scientific formats compatible with common analysis tools and generates visualizations of spatiotemporal wave evolution. By combining numerical accuracy with practical accessibility on modest computational resources, sangkuriang offers a platform suitable for both classroom demonstrations of nonlinear wave phenomena and exploratory research into soliton dynamics.
1Applied Geology Research Group, Bandung Institute of Technology, Bandung 40132, Indonesia
@@
5Applied and Environmental Oceanography Research Group, Bandung Institute of Technology, Bandung 40132, Indonesia
e-mail: dasaptaerwin@itb.ac.id
Keywords: adaptive time integration, Korteweg–de Vries equation, pseudo-spectral methods, soliton dynamics
@@ 1 Introduction
The Korteweg–de Vries (KdV) equation occupies a distinguished position in the undergraduate and graduate physics curriculum, serving as the canonical example of a nonlinear partial differential equation (PDE) that admits exact analytical solutions with remarkable physical properties [3]. First derived in the context of shallow water waves [1], this equation demonstrates how the competing effects of nonlinearity and dispersion can achieve a delicate balance, producing localized wave structures—solitons—that propagate without change of form and survive mutual collisions intact [2]. The discovery that such behavior arises from complete integrability, with an infinite hierarchy of conservation laws accessible through inverse scattering techniques [4, 5], has made the KdV equation a cornerstone for teaching nonlinear dynamics, Hamiltonian systems, and mathematical physics at the advanced undergraduate and graduate levels [6].
Despite the rich analytical structure of soliton theory, students often struggle to develop physical intuition for nonlinear wave phenomena from analytical treatments alone. The interplay between steepening and dispersion, the amplitude-dependent velocity of solitons, and the elastic nature of their collisions are concepts that benefit enormously from direct visualization and numerical experimentation [7]. Computational laboratories, in which students can modify initial conditions, vary physical parameters, and observe the resulting dynamics, offer pedagogical opportunities that complement traditional analytical instruction [8]. Such exercises allow learners to discover the defining characteristics of solitons empirically—for instance, that taller waves travel faster, or that collisions produce phase shifts but preserve wave identities—before or alongside formal derivations.
The choice of programming environment significantly influences the accessibility and effectiveness of computational physics education. Python has emerged as the dominant language for scientific computing instruction, owing to its readable syntax, extensive ecosystem of numerical libraries, and low barrier to entry for students without prior programming experience [9, 10]. The availability of high-quality implementations of fundamental algorithms—fast Fourier transforms (FFT), adaptive differential equation solvers, and array operations—within freely distributed libraries such as NumPy and SciPy [11, 12] enables students to focus on physical concepts rather than low-level implementation details. Furthermore, recent developments in just-in-time (JIT) compilation through libraries such as Numba [13] permit Python codes to achieve performance competitive with compiled languages, removing the traditional trade-off between ease of development and computational speed [14]. These characteristics have established Python as the preferred platform for computational physics courses at institutions worldwide [15, 16].
Pseudo-spectral methods represent a particularly instructive approach to the numerical solution of PDEs, combining the elegance of Fourier analysis with practical computational efficiency [17, 18]. By representing spatial derivatives as multiplications in wavenumber space, these methods achieve high accuracy for smooth periodic solutions while introducing students to concepts—the discrete Fourier transform (DFT), aliasing, spectral convergence—that recur throughout computational physics and signal processing [19, 20]. The KdV equation, with its combination of nonlinear and dispersive terms operating on smooth soliton profiles, provides an ideal context for introducing pseudo-spectral techniques at the graduate level.
@@
This article presents sangkuriang, an open-source Python library for solving the KdV equation, designed with both pedagogical clarity and computational efficiency as objectives. The implementation employs Fourier pseudo-spectral spatial discretization coupled with adaptive eighth-order Runge–Kutta time integration via the Dormand–Prince 8(5,3) (DOP853) method [21, 22]. Performance-critical routines are accelerated through Numba JIT compilation with support for multi-core parallelization, enabling research-grade simulations on modest hardware. The library provides a command-line interface for immediate use with predefined test cases, a Python Application Programming Interface (API) for programmatic access, and outputs simulation data in Network Common Data Format (NetCDF) format following Climate and Forecast (CF) conventions [23] for compatibility with standard analysis tools. Beyond basic simulation capabilities, the solver incorporates diagnostic routines for monitoring conservation laws and generating animated visualizations, enabling students to verify numerical accuracy and explore the mathematical structure of integrable systems. The software is freely available from the Python Package Index (PyPI) and distributed under a permissive open-source license.
2 Methods
@@ 2.1 Model Description
The KdV equation constitutes one of the most fundamental nonlinear PDEs in mathematical physics, governing the propagation of weakly nonlinear, weakly dispersive waves in various physical contexts [1, 2]. We derive this equation from first principles beginning with the three-dimensional Navier–Stokes equations, following the established asymptotic analysis framework [7, 6].
Consider an inviscid, incompressible fluid occupying a domain bounded below by a rigid horizontal bottom at $z=0$ and above by a free surface at $z=h_{0}+\eta(x,y,t)$, where $h_{0}$ denotes the undisturbed depth and $\eta$ represents the free surface displacement (Figure 1). The fluid is subject to gravitational acceleration $g$ acting in the negative $z$-direction.
Figure 1: Schematic of the shallow water wave domain. The fluid occupies a region bounded below by a rigid horizontal bottom at $z=0$ and above by a free surface at $z=h_{0}+\eta(x,t)$, where $h_{0}$ denotes the undisturbed water depth and $\eta(x,t)$ represents the free surface displacement. Gravitational acceleration $g$ acts in the negative $z$-direction, and wave propagation occurs along the $x$-axis with phase speed $c$.
@@
where $\rho$ is the constant fluid density. These equations are supplemented by boundary conditions at the bottom and free surface [24, 25].
At the rigid bottom $z=0$, the impermeability condition requires
@@
We now restrict attention to unidirectional wave propagation along the $x$-axis, setting $v=0$ and $\partial/\partial y=0$. The governing equations (1)–(4) reduce to the two-dimensional system
@@
To derive the shallow water equations, we introduce characteristic scales: $\lambda$ for horizontal length, $h_{0}$ for vertical length, $a$ for wave amplitude, $\lambda/c_{0}$ for time (where $c_{0}=\sqrt{gh_{0}}$ is the linear long-wave speed), $c_{0}$ for horizontal velocity, and $\rho gh_{0}$ for pressure. The dimensionless variables are defined as
@@
Two fundamental small parameters emerge naturally: the amplitude parameter $\alpha=a/h_{0}\ll 1$ measuring wave nonlinearity, and the dispersion parameter $\beta=(h_{0}/\lambda)^{2}\ll 1$ quantifying the shallowness of the water [26].
@@
The factor $\beta$ multiplying the left-hand side of equation (14) indicates that vertical accelerations are negligible in the shallow water limit $\beta\to 0$ [27].
@@
These equations are exact within the hydrostatic shallow water approximation but remain coupled and nonlinear.
To incorporate weak dispersion and derive the KdV equation, we retain terms of order $\beta$ in the vertical momentum equation (14). The regime $\alpha\sim\beta\sim\mathcal{O}(\delta)\ll 1$ with $\delta$ representing a single small parameter defines the Boussinesq scaling, wherein nonlinear and dispersive effects balance to produce solitary wave solutions [26]. Under this scaling, we seek asymptotic solutions in the form of perturbation expansions
@@
where $c_{0}=\sqrt{gh_{0}}$ denotes the linear shallow water wave speed. This formulation makes explicit the physical origins of each term: the second term represents linear wave propagation, the third term captures amplitude-dependent steepening (nonlinearity), and the fourth term describes frequency-dependent dispersion.
@@
where we have relabeled $\chi\to x$ for notational simplicity. In equation (29), $u(x,t)$ represents the wave amplitude field [m], with spatial coordinate $x$ [m] and temporal coordinate $t$ [s]. The remarkable property of this equation lies in the exact balance between the quadratic nonlinear term $\varepsilon u\partial_{x}u$, which steepens wave profiles and would induce shock formation in isolation, and the third-order dispersive term $\mu\partial_{x}^{3}u$, which spreads energy and favors shorter wavelengths. When these mechanisms achieve equilibrium, the equation admits solitary wave solutions—solitons—that propagate without change of form [2].
@@
represents the total wave amplitude integrated over space. The momentum functional
@@
represents the Hamiltonian of the system, with the cubic term $\frac{\varepsilon}{2}u^{3}$ encoding nonlinear potential energy and the quadratic gradient term $\frac{3\mu}{2}(\partial_{x}u)^{2}$ representing dispersive kinetic energy. These invariants (30)–(32) satisfy $dM/dt=0$, $dP/dt=0$, and $dE/dt=0$ along solutions of equation (29), providing essential diagnostics for numerical accuracy [6].
@@
Equation (34) reveals the fundamental nonlinear dispersion relation: taller solitons travel faster, enabling overtaking collisions. The width parameter $w=\sqrt{12\mu/(\varepsilon A)}$ [m] characterizes the spatial extent, demonstrating that larger amplitudes produce narrower, faster solitons. The remarkable stability of these solutions under mutual collisions—emerging unchanged except for phase shifts—constitutes the defining characteristic of soliton behavior and validates the KdV equation as an integrable system [2, 5].
@@ 2.2 Numerical Implementation
The numerical solution of the KdV equation (29) requires spatial discretization of the differential operators and temporal integration of the resulting semi-discrete system. The sangkuriang solver employs the Fourier pseudo-spectral method for spatial derivatives [18, 19], which may provide exponential convergence rates for smooth periodic solutions, combined with adaptive high-order explicit time integration. This approach has been demonstrated to be effective for nonlinear dispersive wave equations [17, 8].
Consider a periodic spatial domain $x\in[x_{\min},x_{\max}]$ with length $L=x_{\max}-x_{\min}$. We discretize this domain using $N$ uniformly spaced grid points $x_{j}=x_{\min}+j\Delta x$ for $j=0,1,\ldots,N-1$, where the spatial resolution is $\Delta x=L/N$. Let $u_{j}(t)=u(x_{j},t)$ denote the semi-discrete solution at grid point $x_{j}$ and time $t$. The periodicity condition $u(x_{\min},t)=u(x_{\max},t)$ is implicitly enforced through the DFT representation. The DFT and its inverse are defined as [29]
@@
The conservation properties described by equations (30)–(32) serve as critical diagnostics for numerical accuracy [33, 8]. For the semi-discrete solution $\mathbf{u}(t)=[u_{0}(t),u_{1}(t),\ldots,u_{N-1}(t)]^{T}$, these integrals are approximated using the composite trapezoidal rule [34]:
@@
The complete numerical algorithm proceeds as follows. Given initial condition $u_{0}(x)=u(x,t=0)$, we sample at grid points to obtain the initial state vector $\mathbf{u}(0)=[u_{0}(x_{0}),u_{0}(x_{1}),\ldots,u_{0}(x_{N-1})]^{T}$. The wavenumber vector $\{k_{n}\}_{n=0}^{N-1}$ is precomputed according to equation (37) and stored for repeated use in spectral derivative evaluations. The Numba JIT compiler is initialized with the specified number of parallel threads via set_num_threads(n_cores), enabling multi-core parallelization for subsequent RHS evaluations. At each time step, the right-hand side operator (41) is evaluated by: (i) computing the FFT of the current state via equation (35), (ii) multiplying by $ik_{n}$ and $(ik_{n})^{3}$ to obtain derivative Fourier coefficients, (iii) applying the IFFT via equation (36) to recover physical space derivatives, and (iv) assembling the nonlinear and dispersive terms using JIT-compiled parallel functions. The DOP853 integrator advances the solution from $t_{n}$ to $t_{n+1}$ according to the adaptive strategy described by equations (43)–(44). At specified output times $\{t_{m}\}_{m=0}^{M}$, the solution state is stored and conservation laws (45)–(47) are evaluated for diagnostic purposes. This process continues until the final time $T_{\text{final}}$ is reached. The sangkuriang numerical implementation leverages established scientific computing libraries: NumPy [11] for array operations and FFT routines, SciPy [12] for adaptive ODE integration, Numba [13] for JIT compilation and parallel processing optimization, Matplotlib [35] for visualization, and netCDF4 [23] for scientific data output following CF-1.8 conventions. This combination of spectral methods, high-order adaptive time integration, and parallel computing optimizations may provide both computational efficiency and numerical accuracy appropriate for research-grade simulations of soliton dynamics.

### v3

# arXiv:2601.12029v3[nlin.PS] 27 Feb 2026

@@ Abstract
The Korteweg–de Vries (KdV) equation is a foundational model in geophysical fluid dynamics (GFD), governing the propagation of long internal and surface gravity waves in stratified and shallow ocean environments where the interplay between nonlinear steepening and frequency-dependent dispersion gives rise to solitons. Although the analytical tractability of the KdV equation through inverse scattering is well established, systematic numerical exploration of multi-soliton interactions remains valuable for benchmarking solvers, probing conservation properties under varied oceanic initial conditions, and building intuition for more complex ocean wave phenomena. This article presents sangkuriang, an open-source Python library that solves the KdV equation using Fourier pseudo-spectral spatial discretization and adaptive eighth-order Runge–Kutta time integration. The implementation leverages just-in-time (JIT) compilation to achieve research-grade computational efficiency on standard hardware. The solver is validated through four progressively complex idealized scenarios motivated by oceanic wave dynamics: isolated soliton propagation, symmetric interactions, overtaking collisions, and three-body interactions. High-fidelity conservation of mass, momentum, and energy is demonstrated, with relative errors remaining below $\mathcal{O}(10^{-4})$ across all test cases. Measured soliton velocities align with theoretical predictions within 5%, confirming the capture of the amplitude-dependent dispersion characteristic of oceanic solitary waves. Complementary diagnostics, including spectral entropy and recurrence quantification analysis (RQA), verify that the numerical solutions preserve the regular phase-space structure characteristic of integrable Hamiltonian systems. These results establish sangkuriang as a robust, lightweight platform for reproducible numerical investigation of idealized nonlinear dispersive wave dynamics relevant to coastal and ocean engineering applications.
1Applied Geology Research Group, Bandung Institute of Technology, Bandung, West Java 40132, Indonesia
@@
5Center for Agrarian Studies, Bandung Institute of Technology, Bandung, West Java 40132, Indonesia
6Applied and Environmental Oceanography Research Group, Bandung Institute of Technology, Bandung, West Java 40132, Indonesia
7Spatial Systems and Cadaster Research Group, Bandung Institute of Technology, Bandung, West Java 40132, Indonesia
8Atmospheric Science Research Group, Bandung Institute of Technology, Bandung, West Java 40132, Indonesia
Correspondence: dasaptaerwin@itb.ac.id
Keywords: Korteweg–de Vries equation; pseudo-spectral methods; soliton dynamics; nonlinear ocean waves; shallow water waves
@@ 1 Introduction
The Korteweg–de Vries (KdV) equation occupies a central position in geophysical fluid dynamics (GFD) as the prototypical completely integrable partial differential equation (PDE) admitting soliton solutions [1]. Originally derived to describe long gravity waves in shallow channels [2], the equation demonstrates how competing nonlinear steepening and linear dispersion can balance to produce localized, shape-preserving pulses—solitons—that survive mutual collisions with only phase shifts [3]. The discovery that this behavior is rooted in complete integrability, with an infinite hierarchy of conservation laws accessible through inverse scattering [4, 5], elevated the KdV equation to a cornerstone of mathematical physics, with applications spanning plasma dynamics, internal ocean waves, and condensed matter systems [6].
In the oceanic context, the KdV framework and its extensions have proven indispensable for describing nonlinear internal solitary waves (ISWs), which are among the most energetic high-frequency phenomena in the coastal and deep ocean [46, 47]. Satellite imagery and in situ observations have documented large-amplitude ISWs in numerous marginal seas, including the South China Sea, the Andaman Sea [49], and the Sulu Sea, where these waves modulate acoustic propagation, drive turbulent mixing, and exert forces on offshore structures [48]. While realistic ocean modeling requires variable-coefficient and higher-order extensions of the KdV equation [48, 50], the constant-coefficient idealization retains its importance as the canonical benchmark for understanding the fundamental balance between nonlinearity and dispersion that underpins oceanic solitary wave dynamics [6, 46].
Although the analytical structure of soliton theory is well established, numerical simulation remains indispensable for exploring parameter regimes, validating approximate initial conditions, and characterizing transient interaction dynamics that resist closed-form description [7]. In particular, the quantitative assessment of conservation law preservation, amplitude-dependent dispersion, and phase-space regularity across multi-soliton configurations requires flexible computational tools capable of high-fidelity integration over extended time intervals [8]. Such capabilities are especially relevant to idealized studies of oceanic internal wave interactions, where multi-soliton collisions and overtaking events are routinely observed in field campaigns [51].
Python has become a widely adopted language for scientific computing in the geosciences and oceanography, owing to its extensive numerical ecosystem and low overhead for prototyping [9, 10]. Libraries such as NumPy and SciPy [11, 12] provide efficient implementations of core algorithms—fast Fourier transforms (FFT), adaptive ODE solvers, and dense linear algebra—while just-in-time (JIT) compilation via Numba [13] bridges the gap between interpreted flexibility and compiled performance [14]. These capabilities have made Python increasingly viable for production-level numerical experiments in computational geophysics [15, 16].
Pseudo-spectral methods are particularly well suited to the KdV equation. By representing spatial derivatives as multiplications in wavenumber space, Fourier-based discretizations achieve exponential convergence for smooth periodic solutions while naturally respecting the spectral structure of dispersive operators [17, 18]. This approach has a long and successful history in the numerical treatment of nonlinear dispersive PDEs arising in ocean wave theory [19, 20].
@@
In this work, we present sangkuriang, an open-source Python library for solving the KdV equation using Fourier pseudo-spectral spatial discretization coupled with adaptive eighth-order Runge–Kutta time integration via the Dormand–Prince 8(5,3) (DOP853) method [21, 22]. Performance-critical routines are accelerated through Numba JIT compilation with multi-core parallelization, enabling research-grade simulations on commodity hardware. The library provides both a command-line interface for rapid deployment with predefined test cases and a Python API for programmatic access. Simulation data are output in Network Common Data Format (NetCDF) following Climate and Forecast (CF) conventions [23], ensuring interoperability with standard oceanographic and geoscientific analysis workflows. The solver additionally incorporates diagnostic routines for monitoring conservation laws, tracking soliton trajectories, and generating animated visualizations. The software is freely available from the Python Package Index (PyPI) under a permissive open-source license.
We revisit the classical benchmarks of KdV soliton dynamics—single-soliton propagation, symmetric two-soliton configurations, overtaking collisions, and three-body interactions—framed as idealized analogs of oceanic solitary wave phenomena, with an emphasis on quantitative diagnostics that go beyond visual inspection. Specifically, we employ information-theoretic measures (spectral entropy, LMC statistical complexity, Fisher information) and recurrence quantification analysis to characterize the phase-space structure of the computed solutions. These complementary diagnostics provide model-independent evidence for the preservation of integrability under numerical integration and offer tools that may prove useful in studying perturbed or extended KdV-type systems relevant to realistic ocean stratification and bathymetric variability, where analytical guarantees are unavailable.
2 Materials and Methods
@@ 2.1 Model Description
The KdV equation constitutes one of the most fundamental nonlinear PDEs in GFD, governing the propagation of weakly nonlinear, weakly dispersive long waves in shallow ocean basins, coastal shelves, and stratified water columns [2, 3]. The derivation presented here proceeds from first principles beginning with the three-dimensional Euler equations for an inviscid ocean layer, following the established asymptotic analysis framework employed in ocean wave theory [7, 6].
Consider an inviscid, incompressible fluid occupying a domain bounded below by a rigid horizontal seabed at $z=0$ and above by a free surface at $z=h_{0}+\eta(x,y,t)$, where $h_{0}$ denotes the undisturbed water depth and $\eta$ represents the free surface displacement (Figure 1). The fluid is subject to gravitational acceleration $g$ acting in the negative $z$-direction. This configuration represents an idealized coastal or continental shelf environment in which long surface gravity waves propagate over a flat bottom—a standard starting point for shallow water wave theory in physical oceanography [24, 46].
Figure 1: Schematic of the shallow ocean wave domain. The fluid occupies a region bounded below by a rigid horizontal seabed at $z=0$ and above by a free surface at $z=h_{0}+\eta(x,t)$, where $h_{0}$ denotes the undisturbed water depth and $\eta(x,t)$ represents the free surface displacement. Gravitational acceleration $g$ acts in the negative $z$-direction, and wave propagation occurs along the $x$-axis with phase speed $c$.
@@
where $\rho$ is the constant fluid density. These equations are supplemented by boundary conditions at the seabed and free surface [24, 25].
At the rigid seabed $z=0$, the impermeability condition requires
@@
We now restrict attention to unidirectional wave propagation along the $x$-axis, setting $v=0$ and $\partial/\partial y=0$. This reduction is appropriate for long-crested ocean waves propagating shoreward over a continental shelf or along a waveguide [47]. The governing equations (1)–(4) reduce to the two-dimensional system
@@
To derive the shallow water equations, we introduce characteristic scales: $\lambda$ for horizontal length (the dominant wavelength), $h_{0}$ for vertical length (the water depth), $a$ for wave amplitude, $\lambda/c_{0}$ for time (where $c_{0}=\sqrt{gh_{0}}$ is the linear long-wave speed), $c_{0}$ for horizontal velocity, and $\rho gh_{0}$ for pressure. The dimensionless variables are defined as
@@
Two fundamental small parameters emerge naturally: the amplitude parameter $\alpha=a/h_{0}\ll 1$ measuring wave nonlinearity, and the dispersion parameter $\beta=(h_{0}/\lambda)^{2}\ll 1$ quantifying the shallowness of the water [26]. In typical continental shelf settings where long gravity waves or internal solitary waves are observed, both $\alpha$ and $\beta$ are indeed small, placing the dynamics squarely within the weakly nonlinear, weakly dispersive regime that the KdV equation governs [46, 48].
@@
The factor $\beta$ multiplying the left-hand side of equation (14) indicates that vertical accelerations are negligible in the shallow water limit $\beta\to 0$, consistent with the hydrostatic approximation widely employed in large-scale ocean circulation models [27].
@@
These equations are exact within the hydrostatic shallow water approximation but remain coupled and nonlinear. They form the basis of many operational coastal and tsunami models, but lack the dispersive corrections necessary to sustain solitary wave solutions [50].
To incorporate weak dispersion and derive the KdV equation, we retain terms of order $\beta$ in the vertical momentum equation (14). The regime $\alpha\sim\beta\sim\mathcal{O}(\delta)\ll 1$ with $\delta$ representing a single small parameter defines the Boussinesq scaling, wherein nonlinear and dispersive effects balance to produce solitary wave solutions [26]. This balance is the essential physical mechanism underlying the formation of internal solitary waves observed in shelf-slope regions worldwide [46, 49]. Under this scaling, we seek asymptotic solutions in the form of perturbation expansions
@@
where $c_{0}=\sqrt{gh_{0}}$ denotes the linear shallow water wave speed. This formulation makes explicit the physical origins of each term: the second term represents linear wave propagation at the long-wave phase speed, the third term captures amplitude-dependent steepening (nonlinearity) responsible for wave breaking in the absence of dispersion, and the fourth term describes frequency-dependent dispersion that spreads energy across wavenumbers. The competition between these last two effects is the fundamental mechanism sustaining solitary waves in shallow ocean environments [46, 48].
@@
where we have relabeled $\chi\to x$ for notational simplicity. In equation (29), $u(x,t)$ represents the wave amplitude field [m], with spatial coordinate $x$ [m] and temporal coordinate $t$ [s]. The remarkable property of this equation lies in the exact balance between the quadratic nonlinear term $\varepsilon u\partial_{x}u$, which steepens wave profiles and would induce shock formation in isolation, and the third-order dispersive term $\mu\partial_{x}^{3}u$, which spreads energy and favors shorter wavelengths. When these mechanisms achieve equilibrium, the equation admits solitary wave solutions—solitons—that propagate without change of form [3]. This idealized balance captures the essential physics of long nonlinear waves observed on continental shelves, in straits, and in the thermocline of the open ocean [47, 51].
@@
represents the total wave amplitude integrated over space, analogous to volume conservation in the ocean surface layer. The momentum functional
@@
represents the Hamiltonian of the system, with the cubic term $\frac{\varepsilon}{2}u^{3}$ encoding nonlinear potential energy and the quadratic gradient term $\frac{3\mu}{2}(\partial_{x}u)^{2}$ representing dispersive kinetic energy. These invariants (30)–(32) satisfy $dM/dt=0$, $dP/dt=0$, and $dE/dt=0$ along solutions of equation (29), providing essential diagnostics for numerical accuracy [6]. Their preservation under numerical integration is a necessary condition for any solver intended to reproduce the long-time dynamics of oceanic solitary waves faithfully.
@@
Equation (34) reveals the fundamental nonlinear dispersion relation: taller solitons travel faster, enabling overtaking collisions. This amplitude–velocity relationship is a robust feature of oceanic internal solitary waves, where larger-amplitude ISWs are consistently observed to propagate faster than their smaller counterparts [46, 49]. The width parameter $w=\sqrt{12\mu/(\varepsilon A)}$ [m] characterizes the spatial extent, demonstrating that larger amplitudes produce narrower, faster solitons—a relationship well documented in field observations of ISWs on continental shelves [48]. The remarkable stability of these solutions under mutual collisions—emerging unchanged except for phase shifts—constitutes the defining characteristic of soliton behavior and validates the KdV equation as an integrable system [3, 5].
@@ 2.2 Numerical Implementation
The numerical solution of the KdV equation (29) requires spatial discretization of the differential operators and temporal integration of the resulting semi-discrete system. The sangkuriang solver employs the Fourier pseudo-spectral method for spatial derivatives [18, 19], which may provide exponential convergence rates for smooth periodic solutions, combined with adaptive high-order explicit time integration. This approach has been demonstrated to be effective for nonlinear dispersive wave equations arising in ocean and atmospheric dynamics [17, 8].
Consider a periodic spatial domain $x\in[x_{\min},x_{\max}]$ with length $L=x_{\max}-x_{\min}$. We discretize this domain using $N$ uniformly spaced grid points $x_{j}=x_{\min}+j\Delta x$ for $j=0,1,\ldots,N-1$, where the spatial resolution is $\Delta x=L/N$. Let $u_{j}(t)=u(x_{j},t)$ denote the semi-discrete solution at grid point $x_{j}$ and time $t$. The periodicity condition $u(x_{\min},t)=u(x_{\max},t)$ is implicitly enforced through the DFT representation. While ocean domains are not strictly periodic, the periodic boundary assumption is standard in idealized KdV studies and is justified when soliton amplitudes decay sufficiently before reaching the domain boundaries [17, 20]. The DFT and its inverse are defined as [29]
@@
The conservation properties described by equations (30)–(32) serve as critical diagnostics for numerical accuracy [33, 8]. In the oceanographic context, the preservation of these invariants ensures that the numerical solver does not introduce spurious sources or sinks of wave energy or momentum that could distort the simulated soliton dynamics. For the semi-discrete solution $\mathbf{u}(t)=[u_{0}(t),u_{1}(t),\ldots,u_{N-1}(t)]^{T}$, these integrals are approximated using the composite trapezoidal rule [34]:
@@
The complete numerical algorithm proceeds as follows. Given initial condition $u_{0}(x)=u(x,t=0)$, we sample at grid points to obtain the initial state vector $\mathbf{u}(0)=[u_{0}(x_{0}),u_{0}(x_{1}),\ldots,u_{0}(x_{N-1})]^{T}$. The wavenumber vector $\{k_{n}\}_{n=0}^{N-1}$ is precomputed according to equation (37) and stored for repeated use in spectral derivative evaluations. The Numba JIT compiler is initialized with the specified number of parallel threads via set_num_threads(n_cores), enabling multi-core parallelization for subsequent RHS evaluations. At each time step, the right-hand side operator (41) is evaluated by: (i) computing the FFT of the current state via equation (35), (ii) multiplying by $ik_{n}$ and $(ik_{n})^{3}$ to obtain derivative Fourier coefficients, (iii) applying the IFFT via equation (36) to recover physical space derivatives, and (iv) assembling the nonlinear and dispersive terms using JIT-compiled parallel functions. The DOP853 integrator advances the solution from $t_{n}$ to $t_{n+1}$ according to the adaptive strategy described by equations (43)–(44). At specified output times $\{t_{m}\}_{m=0}^{M}$, the solution state is stored and conservation laws (45)–(47) are evaluated for diagnostic purposes. This process continues until the final time $T_{\text{final}}$ is reached. The sangkuriang numerical implementation leverages established scientific computing libraries widely used in the oceanographic and geoscience communities: NumPy [11] for array operations and FFT routines, SciPy [12] for adaptive ODE integration, Numba [13] for JIT compilation and parallel processing optimization, Matplotlib [35] for visualization, and netCDF4 [23] for scientific data output following CF-1.8 conventions. This combination of spectral methods, high-order adaptive time integration, and parallel computing optimizations may provide both computational efficiency and numerical accuracy appropriate for research-grade simulations of idealized oceanic soliton dynamics.